OpenStudy (anonymous):

For what value of x satisfies this equation: x^2 -4x = -4?

OpenStudy (anonymous):

would you like working with that answer is -4, btw

OpenStudy (anonymous):

not -4...

OpenStudy (lgbasallote):

you can use factoring to find x.. \[x^2 - 4x + 4 = 0\]

OpenStudy (anonymous):

sorry, -2

OpenStudy (anonymous):

\[x^2 -4x + 4 = 0\] Move the 4 to the right side first. \[(x-2)(x-2)\] Factor it out. \[x =2\] Therefore, x = 2

OpenStudy (anonymous):

great